# ChatSpamDetector: Leveraging Large Language Models for Effective Phishing Email Detection

### v1

@@ 5.5 LLM Capabilities for Phishing Email Detection
Combining Multiple Clues to Identify Evidence LLMs can identify evidence leading to the determination of phishing emails by combining clues extracted from both headers and the body. Equipped with knowledge of brands and their associated legitimate domain names, LLMs can distinguish mismatches between these and the sender’s address or URLs in the email body. Additionally, LLMs accurately recognize the use of domain names similar to legitimate ones or URL shortening services (e.g., tinyurl[.]com and cutt[.]ly) to redirect users to phishing sites, with the intention of evading spam filters. For example, regarding a phishing email with the sender BÍôčkchaiň.čom <notify@gotgel[.]org>, GPT-4 identified it as phishing under the Blockchain.com brand and correctly identified the tactic, stating, “The sender’s use of broken language and special characters to replace normal alphabets shows attempts at bypassing spam filters.” Furthermore, GPT-4 detected 124 phishing emails without impersonating brands, through the presence of SE techniques.
Generating Analysis Reports Our system enables LLMs to output both detailed rationales and simplified explanations, allowing users to understand why an analyzed email is suspected of being phishing or is considered legitimate. This capability supports users in making informed decisions about emails diverted to the spam folder, determining whether to ignore them as phishing attempts or to reconsider them as false positives.
@@ 5.6 False Positive and Negative
False Positive We analyzed false positives, where our system incorrectly identified legitimate emails as phishing. Our system using GPT-4 produced only three false positives. One case involved a legitimate email from Tesco where the domain name of the sender address was different from the legitimate domain name and there were multiple subdomains, leading to misclassification. Another case involved a legitimate email from CNET where the complexity of the local part of the sender’s address led to misclassification, compounded by the presence of multiple links in the email body, which raised suspicion. The third case was a promotional email for a new publication issue; although our system did not identify the brand, phrases such as "GET YOUR FREE NECKLACE NOW" led to misclassification as phishing because they appeared to be SE techniques. On the other hand, GPT-3.5 had 32 false positives. These included misinterpreting SE techniques, failing to determine the legitimacy of the sender’s address, and classifying phishing without evidence. Misinterpretations of SE techniques included responding to words such as “free” or “discount” and sensational content in news articles, such as corporate scandals. GPT-3.5 considered these to be common tactics used to attract users. In contrast, GPT-4 not only focuses on such emotional phrases but also performs a deep analysis to determine the correct SE technique based on how it leads to the phishing site. GPT-3.5 also failed to correctly identify the legitimacy of the sender’s address, for example, misclassifying an email from a user’s Hotmail address as Hotmail phishing. In addition, there were cases where there was no evidence of phishing in the headers or body, yet the final classification was phishing. GPT-4 provided more detailed descriptions of the reasons compared to GPT-3.5. While GPT-4 detailed specific SE techniques and suspicious indicators in the headers, GPT-3.5 often made general statements about the presence of phishing indicators in the body or hyperlinks, or the presence of brand impersonation, without providing concrete evidence.
@@
False Negative We present cases in which the proposed system erroneously classified a phishing email as legitimate. Our system using GPT-4 had only three false negatives. Two of these were phishing emails claiming to be from FTX, both attempting to trick recipients into submitting customer claims by a certain deadline. The system missed the phishing email because the authentication of the domain name was successful and the link in the body contained a legitimate FTX domain name. In another case, an email from Tether Cashback asked the recipient to click a link to activate a new account. The system determined that this was not phishing because it found no spoofing or false information in the email. In comparison, GPT-3.5 had 30 false negatives. In some cases, no phishing evidence was detected at all, while in others, some evidence was detected but the email was misclassified as legitimate when combined with other evidence. As an example of the former, a phishing site posed as Microsoft Privacy, requesting an “update to the latest User Agreement.” It exploited Bing’s redirect feature with a URL starting with https://www.bing[.]com/ck/a?!&&p=, leading to a phishing site. GPT-3.5 correctly recognized the Microsoft brand but mistakenly deemed it legitimate due to the recipient’s server being outlook[.]com, despite the sender’s address being unrelated to Microsoft. On the other hand, GPT-4 correctly identified the phishing email by highlighting the SE techniques of urging users, along with the sender’s address domain name discrepancy. It stated, “The link provided in the email is obfuscated, potentially leading to a malicious site, as it does not direct to a Microsoft domain, instead using bing[.]com”. One notable case involves a phishing email that abused the Google Document sharing feature and was sent from Google Notifications <notify-noreply@google[.]com>. GPT-3.5 detected nothing suspicious in the sender or body, as both were legitimate. However, we confirmed that accessing the Google Form link in the email led to a BTC withdrawal scam site. GPT-4 identified that the content and subject of the email used SE techniques to capture the user’s interest. Surprisingly, it correctly recognized the email as phishing by identifying the common attacker tactic of using Google Forms to direct users to phishing sites. This ability to correctly identify phishing, even in emails that use legitimate services and contain links with legitimate domain names, is an example of GPT-4’s advanced analysis capabilities. Another example is an email with the subject line “Update:Your wallet has failed the Merge” pretending to be from the MetaMask cryptocurrency software wallet. GPT-3.5 noted that the sender address did not belong to the legitimate MetaMask domain name and that the body of the email contained suspicious elements urging the user to take action, but concluded with “Further investigation is required,” thus avoiding a definitive phishing judgment. In contrast, GPT-4 correctly identified these tactics and determined that the email was phishing.
@@ 6 Related Work
Machine Learning and Deep Learning Approaches A wide range of research has been conducted on the detection of phishing emails, including proposals for effective detection methods, analysis of current security features, introduction of new protocols, and exploration of the human factors involved in phishing and their countermeasures. One approach in detecting phishing emails is using machine learning-based techniques, which involve extracting features from email headers and bodies and classifying them using various classifiers [35, 36, 45]. Additionally, deep learning-based natural language processing models have been used to interpret context and detect phishing tendencies. For instance, Li et al.[31] employed Long Short-Term Memory (LSTM) networks for phishing detection in large-scale email data. Che et al.[12] used a combination of semantic web databases and fuzzy logic control in their approach. Several studies have proposed using BERT to analyze context for phishing email detection [30, 38, 40]. Our system aims to conduct more advanced analyses by using LLMs to analyze the context of emails, allowing for a more comprehensive analysis of SE techniques present in both the body and headers.
@@ 7 Discussion
LLM Parameters In our research, we used the default parameters for each LLM, but adjusting these parameters could change the results. We found that the detection accuracy of phishing emails was sufficiently high with the default parameters, and the rationale was thoroughly and concretely explained. However, if consistent answers are necessary or more flexible decision-making is desired, customizing the parameters could allow for adjustments tailored to specific needs.
Improving Response Accuracy with RAG LLMs can enhance their outputs through Retrieval-Augmented Generation (RAG) by searching for information from external knowledge bases. This study analyzed the performance of systems solely using LLMs; however, by employing RAG to retrieve lists of legitimate domain names associated with brands, there is potential to further improve accuracy, particularly in distinguishing legitimate emails.
@@ 8 Conclusion
In this paper, we proposed ChatSpamDetector, a novel system for detecting phishing emails. This system examines both the headers and the body of emails to identify various deceptive strategies, including brand impersonation and social engineering tactics. Moreover, ChatSpamDetector can provide detailed explanations for its determinations, drawing on specific evidence to confirm an email as a phishing attempt. Our evaluation experiments demonstrated that ChatSpamDetector significantly outperforms existing baseline systems, achieving a remarkable detection accuracy of 99.70% by using GPT-4. Unlike existing spam filters that rely on continuous updates to their models and block lists—after analyzing a large corpus of phishing emails—our system excels at identifying a wide range of phishing emails across multiple languages with high accuracy, without necessitating further training. This system not only provides a new option for spam prevention but also enables users to make informed decisions by providing concrete rationales for the suspiciousness of emails.

### v2

@@ 5.5.2 Identifying SE Techniques in the Body
5.5.3 Combining Multiple Clues to Identify Evidence
LLMs can identify evidence leading to the determination of phishing emails by combining clues extracted from both headers and the body. Equipped with knowledge of brands and their associated legitimate domain names, LLMs can distinguish mismatches between these and the sender’s address or URLs in the email body. Additionally, LLMs accurately recognize the use of domain names similar to legitimate ones or URL shortening services (e.g., tinyurl[.]com and cutt[.]ly) to redirect users to phishing sites, with the intention of evading spam filters. For example, regarding a phishing email with the sender BÍôčkchaiň.čom <notify@gotgel[.]org>, GPT-4 identified it as phishing under the Blockchain.com brand and correctly identified the tactic, stating, “The sender’s use of broken language and special characters to replace normal alphabets shows attempts at bypassing spam filters.” Furthermore, GPT-4 detected 124 phishing emails without impersonating brands, through the presence of SE techniques.
5.5.4 Generating Analysis Reports
Our system enables LLMs to output both detailed rationales and simplified explanations, allowing users to understand why an analyzed email is suspected of being phishing or is considered legitimate. This capability supports users in making informed decisions about emails diverted to the spam folder, determining whether to ignore them as phishing attempts or to reconsider them as false positives.
@@ 5.6.1 False Positive
We analyzed false positives, where our system incorrectly identified legitimate emails as phishing. Our system using GPT-4 produced only three false positives. One case involved a legitimate email from Tesco where the domain name of the sender address was different from the legitimate domain name and there were multiple subdomains, leading to misclassification. Another case involved a legitimate email from CNET where the complexity of the local part of the sender’s address led to misclassification, compounded by the presence of multiple links in the email body, which raised suspicion. The third case was a promotional email for a new publication issue; although our system did not identify the brand, phrases such as "GET YOUR FREE NECKLACE NOW" led to misclassification as phishing because they appeared to be SE techniques.
On the other hand, GPT-3.5 had 32 false positives. These included misinterpreting SE techniques, failing to determine the legitimacy of the sender’s address, and classifying phishing without evidence. Misinterpretations of SE techniques included responding to words such as “free” or “discount” and sensational content in news articles, such as corporate scandals. GPT-3.5 considered these to be common tactics used to attract users. In contrast, GPT-4 not only focuses on such emotional phrases but also performs a deep analysis to determine the correct SE technique based on how it leads to the phishing site. GPT-3.5 also failed to correctly identify the legitimacy of the sender’s address, for example, misclassifying an email from a user’s Hotmail address as Hotmail phishing. In addition, there were cases where there was no evidence of phishing in the headers or body, yet the final classification was phishing. GPT-4 provided more detailed descriptions of the reasons compared to GPT-3.5. While GPT-4 detailed specific SE techniques and suspicious indicators in the headers, GPT-3.5 often made general statements about the presence of phishing indicators in the body or hyperlinks, or the presence of brand impersonation, without providing concrete evidence.
5.6.2 False Negative
We present cases in which the proposed system erroneously classified a phishing email as legitimate. Our system using GPT-4 had only three false negatives. Two of these were phishing emails claiming to be from FTX, both attempting to trick recipients into submitting customer claims by a certain deadline. The system missed the phishing email because the authentication of the domain name was successful and the link in the body contained a legitimate FTX domain name. In another case, an email from Tether Cashback asked the recipient to click a link to activate a new account. The system determined that this was not phishing because it found no spoofing or false information in the email.
In comparison, GPT-3.5 had 30 false negatives. In some cases, no phishing evidence was detected at all, while in others, some evidence was detected but the email was misclassified as legitimate when combined with other evidence. As an example of the former, a phishing site posed as Microsoft Privacy, requesting an “update to the latest User Agreement.” It exploited Bing’s redirect feature with a URL starting with https://www.bing[.]com/ck/a?!&&p=, leading to a phishing site. GPT-3.5 correctly recognized the Microsoft brand but mistakenly deemed it legitimate due to the recipient’s server being outlook[.]com, despite the sender’s address being unrelated to Microsoft. On the other hand, GPT-4 correctly identified the phishing email by highlighting the SE techniques of urging users, along with the sender’s address domain name discrepancy. It stated, “The link provided in the email is obfuscated, potentially leading to a malicious site, as it does not direct to a Microsoft domain, instead using bing[.]com”.
@@ 5.6.2 False Negative
One notable case involves a phishing email that abused the Google Document sharing feature and was sent from GoogleNotifications<notify-noreply@google[.]com>. GPT-3.5 detected nothing suspicious in the sender or body, as both were legitimate. However, we confirmed that accessing the Google Form link in the email led to a BTC withdrawal scam site. GPT-4 identified that the content and subject of the email used SE techniques to capture the user’s interest. Surprisingly, it correctly recognized the email as phishing by identifying the common attacker tactic of using Google Forms to direct users to phishing sites. This ability to correctly identify phishing, even in emails that use legitimate services and contain links with legitimate domain names, is an example of GPT-4’s advanced analysis capabilities. Another example is an email with the subject line “Update:Your wallet has failed the Merge” pretending to be from the MetaMask cryptocurrency software wallet. GPT-3.5 noted that the sender address did not belong to the legitimate MetaMask domain name and that the body of the email contained suspicious elements urging the user to take action, but concluded with “Further investigation is required,” thus avoiding a definitive phishing judgment. In contrast, GPT-4 correctly identified these tactics and determined that the email was phishing.
@@ 6 Related Work
Machine Learning and Deep Learning Approaches A wide range of research has been conducted on the detection of phishing emails, including proposals for effective detection methods, analysis of current security features, introduction of new protocols, and exploration of the human factors involved in phishing and their countermeasures. One approach in detecting phishing emails is using machine learning-based techniques, which involve extracting features from email headers and bodies and classifying them using various classifiers [37, 38, 48]. Additionally, deep learning-based natural language processing models have been used to interpret context and detect phishing tendencies. For instance, Li et al. [33] employed Long Short-Term Memory (LSTM) networks for phishing detection in large-scale email data. Che et al. [13] used a combination of semantic web databases and fuzzy logic control in their approach. Several studies have proposed using BERT to analyze context for phishing email detection [32, 40, 42]. Our system aims to conduct more advanced analyses by using LLMs to analyze the context of emails, allowing for a more comprehensive analysis of SE techniques present in both the body and headers. Heiding et al. [22] explored the use of LLMs for generating and detecting the intention behind phishing emails. They compared the success rates of LLM-generated phishing emails with those created by human experts and evaluate the ability of LLMs to detect malicious intent, highlighting both the promising results and potential for misuse. While their study focused on detecting malicious intent, our research provides a detailed analysis of the phishing email detection capabilities of LLMs by creating a dataset containing both phishing emails and legitimate emails. Chataut et al. [12] evaluated the ability of LLMs to identify phishing attempts in email bodies. Si et al. [46] developed a Chinese phishing email dataset and proposed an in-context learning approach for LLMs to classify phishing emails. Their research showed that increasing the number of few-shot examples improved classification accuracy. While these studies primarily focused on analyzing email bodies to identify phishing attempts, our system takes a more comprehensive approach. We use LLMs to analyze all email components, including headers. This enables the detection of various phishing techniques, such as sender spoofing, and provides a more robust identification system.
@@ 7 Discussion
LLM Parameters In our research, we used the default parameters for each LLM, but adjusting these parameters could change the results. For instance, Temperature and Top_p are known to significantly influence the output. Temperature is a parameter that controls the randomness of the generated text, with higher values causing the model to produce more diverse and unpredictable text. Similarly, Top_p is a parameter that regulates the diversity of the generated text, where higher values prompt the model to select from a broader range of high-probability tokens in the probability distribution. We found that the detection accuracy of phishing emails was sufficiently high with the default parameters, and the rationale was thoroughly and concretely explained. However, if consistent answers are necessary or more flexible decision-making is desired, customizing the parameters could allow for adjustments tailored to specific needs.
Improving Response Accuracy with RAG LLMs can enhance their outputs through Retrieval-Augmented Generation (RAG) by searching for information from external knowledge bases. In this study, we attempted to detect phishing emails using only the capabilities of LLMs without using such techniques, and we achieved high accuracy in distinguishing between phishing emails and legitimate emails. This is because the LLMs we tested successfully identified the association between brands targeted by phishing emails and their legitimate domain names through pre-training on a variety of textual data. Even for local brands in non-English regions, GPT-4 in particular was able to identify brand impersonation by comparing legitimate domain names. However, LLMs cannot identify the existence of new brands that have emerged after the training period or newly acquired legitimate domain names. This is where RAG may work effectively to overcome such limitations of LLMs. For example, we can provide additional domain name knowledge to LLMs by creating a list of legitimate domain names or by integrating LLMs with search engines to provide real-time access to up-to-date domain information. Such an extension of the capabilities of LLMs using RAG could be a future research topic.
Deployment Scenarios We discuss the real-world deployment scenarios for our system. Our evaluation experiments have shown that our system has high detection capabilities for multilingual phishing emails impersonating various brands. For real-world deployment, the system could potentially replace existing phishing email defense mechanisms implemented in email services or analyze individual emails on user devices. However, as shown in section 5.2, there is a cost associated with each email analysis. This means that analyzing all emails received by a company or organization using our system could potentially incur significant costs. However, given the intense competition in LLM development, there is an ongoing trend of more powerful model APIs becoming available at lower prices. In fact, as of May 2024, the API costs for gpt-4-turbo have been significantly reduced compared to the gpt-4-0613 model we used in our experiments, with input costs decreasing from \$0.03/1K tokens to \$0.01/1K tokens and output costs from \$0.06/1K tokens to \$0.03/1K tokens. Furthermore, the development of high-performance open-source models is actively progressing, and in the future, the cost of analyzing large volumes of emails may no longer be a concern. Realistic scenarios include using the proposed system for email analysis by security vendors to generate intelligence, or by end users to analyze suspicious emails that have been classified into spam folders or have reached their inbox, on a case-by-case basis. Moreover, using lower-cost models like GPT-3.5, which achieved a high accuracy of 96.92% in our experiments, for the initial analysis and then performing detailed analysis using more accurate models only for results with ambiguous scores could also be an effective approach.
@@ 8 Conclusion
In this paper, we proposed ChatSpamDetector, a novel system for detecting phishing emails. This system examines both the headers and the body of emails to identify various deceptive strategies, including brand impersonation and social engineering tactics. Moreover, ChatSpamDetector can provide detailed explanations for its determinations, drawing on specific evidence to confirm an email as a phishing attempt.
Our evaluation experiments demonstrated that ChatSpamDetector significantly outperforms existing baseline systems, achieving a remarkable detection accuracy of 99.70% by using GPT-4. Unlike existing spam filters that rely on continuous updates to their models and block lists—after analyzing a large corpus of phishing emails—our system excels at identifying a wide range of phishing emails across multiple languages with high accuracy, without necessitating further training. This system not only provides a new option for preventing phishing emails, but also enables users to make informed decisions by providing concrete rationales for the suspiciousness of emails.